By a reduction of 30% a number becomes 210. To make the number equal to 450, the original number is to be increased by
Answer by dkppathak(439)

30% reduce make 210
let original number is x
30 % reduce means 100-30 =70
70% of x= 210
70/100 of x= 210
x=210x100/70 = 300
new number 450 and original is 300
increased value =450-300=150
x/100 of 300 =150
x=50 %
answer 50%